OpenStudy (anonymous):

How do I set this up? The initial value of a car is 25,000. After one year, the value of the car is 21,250. Write an exponential function to model the expected value of the car. Estimate the value of the car after 5 years.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

What is the change in price from 25,000 to 21,250

OpenStudy (anonymous):

3,750

jimthompson5910 (jim_thompson5910):

divide that by the initial price to get ?

OpenStudy (anonymous):

6 2/3

OpenStudy (anonymous):

y(t) = Y(0) k^t 21250 = 25000 k^1 k = 0.85 y(t) = 25000 (0.85)^t y(5) = 25000(0.85)^5 = 11093 value is declining 15% per year (k=0.85)

OpenStudy (anonymous):

got 0.85 from the ratio of the price after one year to the initial price, did this because "exponential function" hint indicates that ratios are key rather than absolute differences. Value is declining 15% per year, same as multiplying by 0.85 year after year.

OpenStudy (anonymous):

Well I understand the 0.85 but not exactly sure about the 15% still

jimthompson5910 (jim_thompson5910):

the 85% or 0.85 comes from 100% - 15% = 85%

OpenStudy (anonymous):

Oh okay.

OpenStudy (anonymous):

Thank you both!

OpenStudy (anonymous):

You are quite welcome.

jimthompson5910 (jim_thompson5910):

btw, you divided in the wrong order and it should be 3750/25000 = 0.15 so that represents a 15% decrease in value